How to find center of mass with respect to carbon atom in carbon monoxide, if the distance of center of hydrogen atom and center of carbon atom is 1.130A.

Grade:11

1 Answers

jitender
6 years ago
If C-H bond length is 1.13AThen CO bond length will be 1.13/3A=0.38ABecause bond order of C-H is 1 and that of CO is 3.Then centre of mass w.r.t C will be[ 12 (0)+16 (0.38)]/12+16= 6.08/28=0.217A